A very reliable baseball player is known to get a hit when at bat 42% of the time.

He is expected to have 102 more times at bat before the end of the season, and he is 40 hits away from breaking a league hitting record.

What is the probability of him breaking that record before the end of the season?

Problem #1:

Round your answer to 4 decimals.
Answer #1

P(hit), p = 0.42

q = 1 - p = 0.58

Number of hits, n = 102

Normal approximation to the binomial, P(X < A) = P(Z < (A - mean)/standard deviation)

Mean = np

= 102 x 0.42

= 42.84

Standard deviation = \sqrt{npq}

= V102 x 0.42 x 0.58

= 4.9847

P(breaking the record) = P(X \geq 40)

= 1 - P(X < 39.5) {with continuity correction}

= 1 - P(Z < (39.5 - 42.84)/4.9847)

= 1 - P(Z < -0.67)

= 1 - 0.2514

= 0.7486